OpenStudy (theviper):

If\[\LARGE{\alpha,\beta}\]are roots of the equation (x-a)(x-b)+k=0 then a, b will be the roots of the equation????? @apoorvk @Eyad @jhonyy9 @jlvm @Hero @AravindG @UnkleRhaukus Plz help:)

OpenStudy (anonymous):

Sum of roots is given by : -B/A Product of roots is given by : C/A

OpenStudy (anonymous):

I think this question is poorly written by the textbook or instructor.

OpenStudy (anonymous):

I believe it's asking if \((x-\alpha)(x-\beta)+k=0\), then, what must \(k=\)? The answer is a bit obvious. Have you thought on this, TheViper?

OpenStudy (theviper):

Is there any typo?? @Limitless

OpenStudy (theviper):

Ques is fully right:)

OpenStudy (anonymous):

I think it will be: \[ab = k - \alpha \beta\] Isn't it @apoorvk ??

OpenStudy (anonymous):

@apoorvk, so, you think this is asking for an equation \(f(x)=0\) in terms of \(\alpha\) and \(\beta\)?

OpenStudy (theviper):

\[\Huge{\color{green}{\cal{Equation:-}}}\]\[\LARGE{\color{blue}{{(x-\alpha)(x-\beta)-k=0??}}}\]

OpenStudy (anonymous):

@TheViper, do you intentionally mean a difference between a and \(\alpha\)? Likewise, b and \(\beta\) are different?

OpenStudy (theviper):

this is answer given in book:)

OpenStudy (apoorvk):

Oh yes I made an error there, thanks @waterineyes So, \(x^2−(a+b)x+(ab−k)=0\) has roots α and β. Now sum of roots = -B/A = (a+b) = α + β ...{i} Product of roots = αβ = ab - k or, ab = k - αβ ...{ii} Now a quadratic with roots 'a' and 'b' is \(x^2−(a+b)x+ab=0\) Substitute the values of 'a+b' and 'ab' from {i} and {ii} respectively, and enjoy!

OpenStudy (anonymous):

It is okay..

OpenStudy (theviper):

can i have the full sol @apoorvk I here also so confused:(

OpenStudy (apoorvk):

Yes @Limitless I agree that the language of the question is pretty misleading. :|

OpenStudy (apoorvk):

@TheViper you're just a step away from the answer, try to ho through this thing just once more pretty slowly, am sure you'll figure out what we're doing in here.

OpenStudy (anonymous):

Just plug in the values of a+b and ab and you will have equation having roots a and b..

OpenStudy (theviper):

how can be sum of roots=-B/A @apoorvk

OpenStudy (anonymous):

It is...

OpenStudy (apoorvk):

@TheViper am sorry I should have mentioned that: Compared with a standard quadratic equation of the form \(Ax^2 + Bx + C =0\) sum of roots = -B/A and product of roots = C/A You must have heard of this I believe?

OpenStudy (theviper):

no

OpenStudy (apoorvk):

No? Umm, what grade are you in currently?

OpenStudy (theviper):

oh I see it with more concentration I know it:) Product of roots=constant term/co-efficient of x^2 Sum of rots=-co-efficient of x/co-efficient of x^2

OpenStudy (apoorvk):

Yes Sire! Now you get it! :) Hmm so does it seem to make any sense now?

OpenStudy (theviper):

k thanx all of u now I can do it:)

OpenStudy (theviper):

yes

OpenStudy (theviper):

thanx @waterineyes @apoorvk :)

OpenStudy (apoorvk):

You're welcome!

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):